OpenStudy (anonymous):

Use the method of substitution to solve the system of linear equations 3a=5-b 3a=4-b

OpenStudy (anonymous):

@thivitaa

OpenStudy (mertsj):

Use the transitive property. If two quantities are equal to the same thing, then they are equal to each other.

OpenStudy (thivitaa):

a=(5-b)/3-->(1) then sub in this..and try figuring the ans...goodluck..:D

OpenStudy (mertsj):

5-b=3a and 3a = 4-b So 5-b=4-b

OpenStudy (thivitaa):

that could work too...:D

OpenStudy (anonymous):

-b+b ?????

OpenStudy (mertsj):

yes. -b+b=0

OpenStudy (anonymous):

-,,-

OpenStudy (anonymous):

would it be considered no solution?